OpenStudy (anonymous):

hello what is the place value for: 1 777

OpenStudy (anonymous):

Whose place vaule you want to find???

OpenStudy (anonymous):

any :)

OpenStudy (anonymous):

For 7 in unit place : 7*1 = 7 is the place value.. For 7 in ten's place = 7*10 = 70 is the place value.. For 7 in hundred's place = 7*100 = 700 is the place value.. For 1 in thousand's place = 1*1000 = 1000 is the place value..